OpenStudy (anonymous):

Evaluate. 8x3+6 ------ -2[6/2-(9-1)

OpenStudy (anonymous):

@iGreen @CausticSyndicalist @Nnesha

OpenStudy (solomonzelman):

sure, lets start from a question interpretation. \(\large\color{black}{ \frac{\LARGE 8\times3+6}{\LARGE -2[\frac{6}{2}-(9-1)~]} }\) is it like this?

OpenStudy (anonymous):

no

OpenStudy (solomonzelman):

\(\large\color{black}{ \frac{\LARGE 8\times3+6}{\LARGE -2[\frac{6}{2-(9-1)}~]} }\) this?

OpenStudy (anonymous):

no its /= divide 6 by 2

OpenStudy (solomonzelman):

can you draw it (Click "draw" below)

OpenStudy (anonymous):



OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

you can use a pencil to draw it. (sorry this makes it not much better to understand)

OpenStudy (anonymous):